Determine the effective capacitance if two capacitors, C1 = 10F and C2 = 5F, are

connected in series.


1
Expert's answer
2022-05-22T14:38:32-0400
"C=\frac{C_1C_2}{C_1+C_2}"

"C=\frac{10*5}{10+5}=3.3\:\rm F"